OpenStudy (anonymous):

What is the slope of the line y = 3

jimthompson5910 (jim_thompson5910):

hint: y = 3 is the same as y = 0x+3

jimthompson5910 (jim_thompson5910):

compare y = 0x+3 to y = mx+b

OpenStudy (anonymous):

So then it would be 3 right?

jimthompson5910 (jim_thompson5910):

another hint: m is the slope of y = mx+b

OpenStudy (anonymous):

Ohh so the slope is 0

jimthompson5910 (jim_thompson5910):

yes m = 0 is the slope

jimthompson5910 (jim_thompson5910):

when the slope is 0, you have a completely horizontal flat line